OpenStudy (anonymous):

Find the exact value of the expression. sin (sin^-1(2/3) + cos^-1(1/3))

OpenStudy (tkhunny):

You may wish to refer to the identity: \(\sin(A+B) = \sin(A)\cos(B)+\cos(A)\sin(B)\)

OpenStudy (kitkat16):

do you have choices

OpenStudy (anonymous):

I know the answer is 2+2sqrt10/9, but i need to know how to do it